# Mind the Gap: Aligning Knowledge Bases with User Needs to Enhance Mental Health Retrieval

### v1

@@ Appendix A List of Main Topics and Subtopics for Categorization
This appendix provides the full set of main topics and verbatim subtopics used in our LLM-based categorization. Main topics are ordered by their frequency of endorsement in the CLICC dataset, i.e., the proportion of clients for whom each concern was recorded as either a top concern or a check-all-that-apply concern [for Collegiate Mental Health, 2019].
@@ Appendix C Ranked Subtopics by Hybrid Gap Score
This appendix lists subtopics with their corresponding main topics and combined hybrid gap scores. Higher scores indicate larger content gaps and thus a greater need to add documents. If the usefulness score is marked as N/A, this means that no documents were available for that subtopic. The hybrid score was calculated as a 50% coverage score and 50% usefulness score when both were available, or as 100% coverage when usefulness was N/A.
@@ Appendix D Therapist-Reviewed LLM Usefulness Scoring Prompt for Document Helpfulness
Appendix E LLM Prompt for Synthetic Document Generation

### v2

@@ Appendix A List of Main Topics and Subtopics for Categorization
This appendix provides the full set of main topics and verbatim subtopics used in our LLM-based categorization. Main topics are ordered by their frequency of endorsement in the CLICC dataset, i.e., the proportion of clients for whom each concern was recorded as either a top concern or a check-all-that-apply concern [9].
@@ Appendix C Ranked Subtopics by Hybrid Gap Score
Table LABEL:tab:ranked_subtopics lists subtopics with their corresponding main topics, coverage scores, usefulness scores, and combined hybrid gap scores. Higher scores (between 1–100) indicate larger content gaps and thus a greater need to add documents. If the usefulness score is marked as N/A, this means that no documents were available for that subtopic. If the coverage score is low but the usefulness score is high (e.g., career transitions and instability), it indicates that although few resources exist on this topic, those available are highly relevant and valuable to user queries. The hybrid score was calculated as a 50% coverage score and 50% usefulness score when both were available, or as 100% coverage when usefulness was N/A.
@@ Appendix D Therapist-Reviewed LLM Usefulness Scoring Prompt for Document Helpfulness
Appendix E Impact of Coverage-Usefulness Weighting on Corpus Allocation
40/60 0.205 3.5
60/40 0.263 4.5
30/70 0.474 8.1
70/30 0.497 8.5
100/0 1.264 21.6
0/100 1.415 24.2
Appendix F LLM Prompt for Synthetic Document Generation
@@
Appendix G Comparative Analysis of Usefulness Scores Between Directed vs. Non-Directed Corpus Configurations Across RAG Pipelines